OpenStudy (anonymous):

Please help Find the specific values of x on the interval [0,2pi] sin(2x)tan(2x)+sin(2x)=0

OpenStudy (anonymous):

I tried converting tan to sin/cos And then simplifying from there by converting everything into sines and multiplying, but I have no idea where to go from here

OpenStudy (anonymous):

Umm I need to do this algebraically so I can't just go through an check them I need to simplify the equation

OpenStudy (mathmale):

I notice that your expression can be factored. Would you please factor it now.

OpenStudy (anonymous):

Oh sorry I was gone Ok so I got sin2x(tan2x+1) Do I expand using addition properties?

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle ab+a=0 }\) \(\color{#000000 }{ \displaystyle (a\times b)+(a\times 1)=0 }\) \(\color{#000000 }{ \displaystyle a(b+1)=0 }\) \(\color{#000000 }{ \displaystyle a=0 }\) or, \(\color{#000000 }{ \displaystyle b+1=0\quad \Longrightarrow\quad b=-1 }\)\(\small .\)

OpenStudy (solomonzelman):

You are just applying the basic algebra, but to trigonometric functions, rather than just numbers.

OpenStudy (anonymous):

Oh I see that Lol I was thinking more complicated Thank you!

OpenStudy (solomonzelman):

Well, what you have at the moment (which is correct) is\(\small : \) \(\color{#0000ff }{ \displaystyle \sin(2x)\left[\tan(2x)+1\right]=0}\) Then, by the "zero property" you have\(\small : \) \(\color{#0000ff }{ \displaystyle \sin(2x)=0\quad}\)or,\(\color{#0000ff }{ \displaystyle \quad \tan(2x)+1=0}\)

OpenStudy (anonymous):

Yeah I got the specific solutions of 0, pi/2,pi,3pi/2,2pi,3pi/8,7pi/8,11pi/8, and 15pi/8 Is that correct?

OpenStudy (solomonzelman):

\(\color{#0000ff }{ \displaystyle \sin(2x)=0\quad {\small \rm then}\quad x=0,~~\frac{\pi}{2},~~\frac{3\pi}{2},~~2\pi}\) \(\color{#0000ff }{ \displaystyle \tan(2x)+1=0\quad {\small \rm then}\quad x=\frac{3\pi}{8},~~\frac{7\pi}{8},~~\frac{11\pi}{8},~~\frac{15\pi}{8}}\)

OpenStudy (solomonzelman):

yes, your answers are all correct.

OpenStudy (solomonzelman):

(I am lagging a little, sorry takes time to type it)

OpenStudy (anonymous):

Alright thank you!

OpenStudy (solomonzelman):

YW

OpenStudy (danjs):

the first one is for values where sin(x) and cos(x) are zero, what is the second for the double angle of the tangent